OpenStudy (jkbo):

Please help What is the solution to the system of equations? -4x+4y-2z=-8 -3x-y+4z=0 2x-2y+3z=-4 a.(1, –7, –2) b.(–3, –7, –4) c.(–3, 0, 5) d.(3, 5, –4)

OpenStudy (zzr0ck3r):

is this for linear algebra class?

OpenStudy (zzr0ck3r):

there are many ways to do this, so what are you doing in class?

OpenStudy (zzr0ck3r):

@jkbo?

OpenStudy (anonymous):

@zzr0ck3r i am working on this chapter

OpenStudy (zzr0ck3r):

?

OpenStudy (zzr0ck3r):

well if its for linear algebra we need to put it in a matric and row reduce, but I don't want to go through that if it is not for linear algebra. If is NOT for L.A then there are a few methods still and again don't want to waste time showing one when we need the other.

OpenStudy (zzr0ck3r):

matrix*

OpenStudy (jkbo):

yes this is alg. 2

OpenStudy (zzr0ck3r):

is that linear algebra?

OpenStudy (jkbo):

no I dont think so

OpenStudy (anonymous):

the answer is b (–3, –7, –4)

OpenStudy (zzr0ck3r):

@jkbo do you want to know how?

OpenStudy (anonymous):

This is how you get -4x+4y-2z=-8 -3x-y+4z=0 X -2 Multiply by -2 to match with third equation 2x-2y+3z=-4 -4x+4y-2z=-8 ______________________ 6x+2y-8z=0 2x-2y+3z=-4 ______________________ 8x -5z =-4 Add second and third equation, this will eliminate y variable from second and third equation You will get 8x -5z =-4 Now you can multiply -3x-y+4z=0 X 4 Multiply by 4 to match with First equation The reason we want to do is to eliminate Y variable from all three equations after multiplication, you need to add equation first and second, this way it will cross out y variable from both first and second equations -4x + 4y - 2z =-8 -12x -4y +16z =0 _______________________ -16x +14z = -8 _______________________ Now you get two equations after eliminating all three equations. Equation1 = 8x -5z = -4 Equation2 = -16x +14z = -8 We don't have y variable now, we need to eliminate x. After eliminating x we will get z value. To eliminate x, we need to multiply first equation with 2 8x -5z = -4 X 2 -16x +14z = -8 You will get 16x - 10z = -8 -16x +14z = -8 ______________________ 4z = -16 So z = -4 ______________________ Now plug z value in the second equation we got, its your choice either first or second equation. I use second equation to get x value. -16x +14z = -8 -16x +14(-4) = -8 -16x - 56 =-8 -16x = -8+56 -16x = 48 x = -3 So you get x and z value. Plug both values in any of the three equations in the question to get the y value. so i use equation 1 -4x+4y-2z=-8 -4(-3)+4y-2(-4) = -8 12+4y+8 = -8 20+4y = -8 4y = -8-20 4y = -28 y = -7 This is how you get the result X = -3 Y = -7 Z = -4